### v1

# Exact State Evolution and Energy Spectrum in Solvable Bosonic Models

Abstract
Solvable bosonic models provide a fundamental framework for describing light propagation in nonlinear media, including optical down-conversion processes that generate squeezed states of light and their higher-order generalizations. In quantum optics a central objective is to determine the time evolution of a given initial state. Exact analytic solution to the state-evolution problem is presented, applicable to a broad class of solvable bosonic models and arbitrary initial states. Moreover, the characteristic equation governing the energy spectrum is derived and the eigenstates are found in the form of continued fractions and as principal minors of the associated Jacobi matrix. The results provide a solid analytical framework for discussion of exactly solvable bosonic models.
I Introduction
For a broad class of bosonic models describing interactions among several bosonic modes, the Hilbert space decomposes into a direct sum of finite-dimensional invariant subspaces. Within each subspace, the Hamiltonian assumes the form of a tridiagonal Hermitian matrix. These two structural features—the decomposition of the Hilbert space into finite-dimensional invariant subspaces and the tridiagonal representation of the Hamiltonian therein—characterize a class of exactly solvable bosonic models whose energy spectra has been studied by the group-theoretic techniques GroupMeth ; LieAlg3Boson ; GroupMeth2 ; LieAlgGen1 ; LieAlgGen2 , by employing the Bethe ansatz SpinSystems ; 3BosonBethe , or by the Quantum Inverse Scattering Method QISM .
In quantum optics, the propagation of optical modes in a lossless nonlinear medium, under the condition of phase matching, can be described by a solvable bosonic model. In a lossless medium, energy conservation is governed by the Manley–Rowe relations NonlOpt , which ensure that the total optical energy remains constant during propagation. Accordingly, the free-propagation term $\hat{H}_{0}$ of the quantum Hamiltonian, representing the optical energy, is quadratic in the bosonic creation and annihilation operators, while the interaction term $\hat{H}$, of higher order in these operators, governs the mode-conversion processes mediated by the nonlinear medium. When the phase-matching condition is satisfied, the interaction term preserves the total optical energy and therefore commutes with $\hat{H}_{0}$, $\hat{H}_{0},\hat{H}]=0$. Prominent examples include $k$-photon down-conversion processes GenSqueez ; 3phDC , which generalize the second-order down-conversion mechanism responsible for the generation of twin photons and squeezed states of light ReviewSq ; Couteau18 ; Zhang21 .
In quantum-optical applications of solvable bosonic models, the central problem is the determination of the time evolution of the system’s quantum state. Under strong pump condition, the so-called parametric approximation is commonly employed, in which the bosonic operators corresponding to the pump mode are replaced by complex scalars. This approximation leads to the well-known description of squeezed states of light ReviewSq . Beyond the parametric approximation, however, the analysis of solvable bosonic models generally requires more advanced algebraic frameworks, such as deformed Lie algebras or the Quantum Inverse Scattering Method, both of which involve substantial mathematical complexity. Consequently, most physical treatments rely on approximate approaches, including WKB-type techniques WKB4GenFun ; Hillery84 ; WKBpump ; Crouch88 and numerical simulations Reid88 ; Drobny92 ; Buzek93 ; Drobny94 ; Hillery95 , often supported by reductions based on conservation laws. Extending the analysis beyond the parametric approximation is of particular importance, since this regime is expected to exhibit pronounced non-Gaussian quantum effects BeyondPA1 ; BeyondPA2 .
Recently, a simple algebraic method was found My1 for obtaining the exact solution of the state-evolution problem in a broad class of solvable bosonic models. The method is particularly applicable to systems arising in quantum-optical contexts, such as $k$-photon down-conversion and related processes, for which the explicit time evolution was derived for a specific initial state in Ref. My1 . The purpose of the present work is to apply that approach to find the evolution of an arbitrary initial state and to derive explicit expressions for the corresponding eigenstates and for the characteristic equation determining the energy spectrum.
The structure of the paper is as follows. In Section II, the class of bosonic models to which the present approach applies is outlined. Section III provides the solution of the state-evolution problem for an arbitrary initial state by evaluating the average of the unitary evolution operator between two arbitrary basis states within each invariant subspace of the Hilbert space. Theorem1 and Corollaries1 and 2 summarize the main results of this section. In Section IV, the characteristic polynomial determining the energy eigenvalues in each invariant subspace is derived, and the quantum amplitudes of the corresponding eigenstates are expressed in terms of continued fractions and as Jacobi determinants. Theorem 2 and Corollary 3 contain the main results concerning the energy spectrum. Furthermore, the stationary state (the eigenstate with zero energy) is obtained in explicit analytic form. Section V discusses the importance of the results and their possible use in applications. Finally, Section VI presents a brief summary of the results and points on an open problem.
II Solvable bosonic models
The solvable bosonic models considered here are characterized by two main features. In the interaction picture, they can be described as follows (for further details, see Ref. My1 ):
(i) Invariant partition of the Hilbert space.—The (infinite-dimensional) Hilbert space $\mathcal{H}$ decomposes into a direct sum of finite-dimensional subspaces, each invariant under the action of the interaction Hamiltonian $\hat{H}$. One can label the invariant subspaces by their dimension $N+1$:
$\mathcal{H}=\bigoplus_{N}\mathcal{H}_{N},\quad\mathcal{H}_{N}=\mathrm{Span}\{|\Psi^{(N)}_{0}\rangle,|\Psi^{(N)}_{1}\rangle,\ldots,|\Psi^{(N)}_{N}\rangle\}$. (1)
Without loss of generality, if multiple subspaces share the same dimension, an additional label, e.g., $\mathcal{I}=(N,\ell)$, can be introduced to distinguish them. In the following, we focus on a single invariant subspace $\mathcal{H}_{N}$, and therefore omit the subspace indices.
(ii) Ladder-type structure of the interaction Hamiltonian.—The interaction Hamiltonian can be expressed as the sum of two Hermitian-conjugate operators,
$\hat{H}=\hat{A}+\hat{A}^{\dagger},$ (2)
which generate transitions between nearest-neighbor basis states within each invariant subspace:
$\langle\Psi_{n}|\hat{A}|\Psi_{m}\rangle=\langle\Psi_{m}|\hat{A}^{\dagger}|\Psi_{n}\rangle=0,\quad m\neq n+1.$ (3)
It is always possible to choose the phases of the basis states in Eq. (1) such that all nonzero matrix elements are non-negative scalars (generally distinct for each invariant subspace). With this choice, the ladder operators act as
$\hat{A}|\Psi_{n+1}\rangle=\sqrt{\beta_{n}}|\Psi_{n}\rangle,\quad\hat{A}^{\dagger}|\Psi_{n}\rangle=\sqrt{\beta_{n}}|\Psi_{n+1}\rangle,$ (4)
where $\beta_{n}\geq 0$ and $\beta_{N}=0$ in subspace $\mathcal{H}_{N}$ due to the finite dimension $\mathrm{dim}\mathcal{H}_{N}=N+1$.
As discussed in Ref.My1 , many bosonic models describing interactions among two or more optical modes in lossless nonlinear media—where the total optical energy of the interacting modes is conserved—possess the two features outlined above. Exactly integrable bosonic models, whose energy spectra have been analyzed in Refs. GroupMeth ; GroupMeth2 ; LieAlg3Boson ; LieAlgGen1 ; LieAlgGen2 , also belong to this class.
For a concrete example, consider a sub-class of nonlinear bosonic models characterized by the ladder operator
$\hat{A}=(\hat{a}^{\dagger})^{m}\prod_{s=1}^{S}\hat{b}_{s}^{k_{s}},$ (5)
where $m\geq 1$, $S\geq 1$ and $k_{s}\geq 1$ are arbitrary integers. In each invariant subspace of Eq. (1) the state $|\Psi_{0}\rangle$ is annihilated by the ladder operator: $\hat{A}|\Psi_{0}\rangle=0$. In the simplest two-mode case ($S=1$) with $\hat{A}=(\hat{a}^{\dagger})^{m}\hat{b}^{k}$, the invariant subspaces are labeled by the composite index $\mathcal{I}\equiv(N,\ell)$, where $0\leq\ell\leq k-1$. In each subspace $\mathcal{H}_{\mathcal{I}}$, the basis states $|\Psi^{(\mathcal{I})}_{n}\rangle$ are given by the Fock states of the two modes:
$|\Psi^{(\mathcal{I})}_{n}\rangle\equiv|M-mn,kn+\ell\rangle,\quad 0\leq n\leq N\equiv\left[\frac{M}{m}\right],$
$|M-mn,kn+\ell\rangle\equiv\frac{(\hat{a}^{\dagger})^{M-mn}(\hat{b}^{\dagger})^{kn+\ell}}{\sqrt{(M-mn)!(kn+\ell)!}}|Vac\rangle,$
where $\ldots$] denotes the integer part. The corresponding $\beta^{(\mathcal{I})}_{n}$ in each invariant subspace $\mathcal{H}_{\mathcal{I}}$ is determined from Eqs. (4) and (II):
$\beta^{(\mathcal{I})}_{n}=\left[\prod_{i=0}^{m-1}(M-mn-i)\right]\prod_{j=1}^{k}(kn+\ell+j).$ (7)
Since $N$ in Eq. (II) is integer, there is an integer $q$, such that $M=Nm+q$ and $0\leq q\leq m-1$. The finite dimension of the subspace, $\mathrm{dim}(\mathcal{H}_{\mathcal{I}})=N+1$, is insured by $\beta^{(\mathcal{I})}_{N}=0$ in Eq. (76), where at least one factor in the first product vanishes.
In the general multi-mode case of Eq. (5), the invariant subspaces are labeled by the composite index $\mathcal{J}\equiv(N,\ell_{1},\ldots,\ell_{S})$, where $0\leq\ell_{s}\leq k_{s}-1$. The corresponding basis states are Fock states, analogous to Eq. (II). In this case, the $\beta$-parameter reads
$\beta^{(\mathcal{J})}_{n}=\left[\prod_{i=0}^{m-1}(M-mn-i)\right]\prod_{s=1}^{S}\prod_{j=1}^{k_{s}}(k_{s}n+\ell_{s}+j).$ (8)
In the optical applications ClasskSq ; GenSqueez one typically has $m=1$ in Eq. (5), where the boson operator $\hat{a}$ describes the pump mode, whereas the optical modes labeled by $s$ correspond to the output (signal) modes. For example, the $k$-photon down-conversion process into a single signal mode GenSqueez is a special case of Eq. (5) with $m=S=1$ and arbitrary $k\equiv k_{1}$. In this case, the interaction Hamiltonian reads
$\hat{H}=\hat{a}^{\dagger}\hat{b}^{k}+\hat{a}(\hat{b}^{\dagger})^{k}$ (9)
and the corresponding ladder coefficients in the invariant subspace are
$\beta^{(N)}_{n}=(N-n)\prod_{j=1}^{k}(kn+j).$ (10)
III Evolution of arbitrary state
We introduce the auxiliary state-number operator $\hat{n}$, which in each invariant subspace $\mathcal{H}_{N}$ acts as
$\hat{n}|\Psi_{n}\rangle=n|\Psi_{n}\rangle,\quad n=0,\ldots,N.$ (11)
The operators $\hat{n}$, $\hat{A}$, and $\hat{A}^{\dagger}$ satisfy the deformed boson algebra:
$\hat{n},\hat{A}^{\dagger}]=\hat{A}^{\dagger},\quad[\hat{n},\hat{A}]=-\hat{A}$, (12)
since
$\hat{A}\hat{A}^{\dagger}=\beta_{\hat{n}},\quad\hat{A}^{\dagger}\hat{A}=\beta_{\hat{n}-1}.$ (13)
Here $\beta_{\hat{n}}$ is obtained by replacing the scalar index “nn” in $\beta_{n}$ by the operator $\hat{n}$ Eq. (11) and we set $\beta_{-1}\equiv 0$. From Eq. (12) one also obtains the identity valid for an arbitrary scalar function $F(x)$:
$F(\hat{n})\hat{A}^{\dagger}=\hat{A}^{\dagger}F(\hat{n}+1).$ (14)
We work in the interaction picture, where the Hamiltonian is given by Eq. (2) (recall that $\hat{H}_{0},\hat{H}]=0$, with $\hat{H}_{0}$ the free Hamiltonian). Consider the evolution of an arbitrary initial state within the subspace $\mathcal{H}_{N}$, using dimensionless variables such as the dimensionless time $\tau$. To this end, we analyze the evolution of each basis state in Eqs. (1)–(4). To simplify the expressions, below we will mainly employ the rescaled states $(-i\hat{A}^{\dagger})^{k}|\Psi_{0}\rangle$, $k=0,\ldots,N$, and expand the unitary evolution operator in this basis as
$e^{-i\tau(\hat{A}+\hat{A}^{\dagger})}(-i\hat{A}^{\dagger})^{k}|\Psi_{0}\rangle=\sum_{n=0}^{N}\gamma_{n,k}(\tau)(-i\hat{A}^{\dagger})^{n}|\Psi_{0}\rangle.$ (15)
The factor $(-i)^{n}$ is introduced to ensure that the expansion coefficients $\gamma_{n,k}(\tau)$ are real-valued functions.
III.1 The states in the expansion of the evolution operator
Expanding the unitary evolution operator $e^{-i\tau(\hat{A}+\hat{A}^{\dagger})}$ on the left-hand side of Eq. (15) generates states of the form
$|A_{m,k}\rangle\equiv(\hat{A}+\hat{A}^{\dagger})^{m}(\hat{A}^{\dagger})^{k}|\Psi_{0}\rangle.$ (16)
As shown in Appendix A, the state in Eq. (16) can be expressed as
$|A_{m,k}\rangle$ $=$ $\sum_{l=0}^{L_{m,k}}a^{(l)}_{m,k}(\hat{A}^{\dagger})^{k+m-2l}|\Psi_{0}\rangle,$ (17)
$L_{m,k}$ $\equiv$ $\mathrm{min}\left(m,\left[\frac{k+m}{2}\right]\right),$
where $\ldots$] denotes the integer part, and $a^{(l)}_{m,k}$ are scalar coefficients, which are combinations of the $\beta_{n}$ parameters defined in Eq. (4).
Let us derive the recursion relation for the scalars $a^{(l)}_{m,k}$. Multiplying Eq. (17) on the left by $\hat{A}+\hat{A}^{\dagger}$ and using Eq. (13) yields
$|A_{m+1,k}\rangle$ $=$ $\left[\sum_{l=0}^{L_{m,k}}a^{(l)}_{m,k}(\hat{A}^{\dagger})^{k+m+1-2l}\right$.
$\left.+\sum_{l=0}^{L_{m,k}}a^{(l)}_{m,k}\beta_{\hat{n}}(\hat{A}^{\dagger})^{k+m-1-2l}\right]|\Psi_{0}\rangle.$
In the second sum, applying Eq. (14) repeatedly with $F(\hat{n})=\beta_{\hat{n}}$ and introducing a new summation index $l^{\prime}=l+1$ gives
$|A_{m+1,k}\rangle=\left[\sum_{l=0}^{L_{m,k}}a^{(l)}_{m,k}(\hat{A}^{\dagger})^{k+m+1-2l}\right$. (18)
$\left.+\sum_{l^{\prime}=1}^{L_{m,k}+1}\beta_{k+m+1-2l^{\prime}}a^{(l^{\prime}-1)}_{m,k}(\hat{A}^{\dagger})^{k+m+1-2l^{\prime}}\right]|\Psi_{0}\rangle.$
Since $\beta_{-1}=0$ (and more generally $\beta_{p}=0$ for $p<0$), we require $k+m+1\geq 2l^{\prime}$. Moreover, by definition in Eq. (17), $a^{(l)}_{m,k}\equiv 0$ for $l>m$. Hence, the upper limit $L_{m,k}+1$ in the second sum in Eq. (18) can be replaced with $L_{m+1,k}$. The same replacement can be performed in the first sum if we postulate
$a^{(l)}_{m,k}\equiv 0,\quad l\notin[0,L_{m,k}].$ (19)
Finally, combining the two sums in Eq. (18) and comparing with Eq. (17) for $m+1$ (then replacing $m+1\to m$) yields the desired recursion relation:
$a^{(l)}_{m,k}=a^{(l)}_{m-1,k}+\beta_{k+m-2l}a^{(l-1)}_{m-1,k}$. (20)
The boundary values of $a^{(l)}_{m,k}$ for $l=0$ and $l=m$ correspond to the simplest cases of the recursion in Eq. (20). Setting $l=0$ in Eq. (20) and taking into account Eq. (19) we obtain
$a^{(0)}_{m,k}=a^{(0)}_{m-1,k}=\ldots=a^{(0)}_{0,k}=1.$ (21)
Setting $l=m$ for $m\leq k$ in Eq. (20) and using $a^{(0)}_{0,k}=1$ we get:
$a^{(m)}_{m,k}=\beta_{k-m}a^{(m-1)}_{m-1,k}=\ldots=\prod_{s=k-m}^{k-1}\beta_{s},\quad\forall m\leq k.$ (22)
For $m>k$, Eq. (22) yields zero, since $\beta_{p}=0$ for $p<0$. This agrees with the condition on $a^{(l)}_{m,k}$ in Eq. (19), which allows non-zero values only for $l\leq L_{m,k}=[(k+m)/2]<m$ in this case.
III.2 Solving the recursion; the $g$-factors
It proves convenient for what follows to reindex the scalars $a^{(l)}_{m,k}$ of Eq. (17) by introducing new ones $g^{(l)}_{n,k}$, defined as
$g^{(l)}_{k+m-2l,k}\equiv a^{(l)}_{m,k}$. (23)
Then, Eq. (19) together with the relation m=nk+2lm=n-k+2l can be cast in the form
$g^{(l)}_{n,k}\equiv 0,\quad l<0\cup n<0\cup 2l<k-n.$ (24)
Let us also rewrite the recursion in Eq. (20) in terms of $g^{(l)}_{n,k}$. We get
$g^{(l)}_{n,k}=g^{(l)}_{n-1,k}+\beta_{n}g^{(l-1)}_{n+1,k}$. (25)
The boundary conditions in Eqs. (21)-(22) become
$g^{(0)}_{n,k}=1,\quad n\geq k,$ (26)
$g^{(k-n)}_{n,k}=\prod_{s=n}^{k-1}\beta_{s},\quad n\leq k-1.$ (27)
If the empty product is set to unity, then the two conditions in Eq. (25) coincide for $n=k$. We will adopt this convention below.
Now, it is easy to show that $g^{(l)}_{n,k}=0$ for l<knl<k-n, i.e., the (nonzero) boundary values of $g^{(l)}_{n,k}$ for $n<k$ are given by Eq. (27). To this goal let us rewrite Eq. (25) for $n+1$ in the form suitable for induction:
$g^{(l)}_{n,k}=g^{(l)}_{n+1,k}-\beta_{n+1}g^{(l-1)}_{n+2,k}$. (28)
For $k-n-1\geq 1$, i.e., $n+1\leq k-1$ by setting l=kn1l=k-n-1 in Eq. (28) and using Eq. (27) we obtain:
$g^{(k-n-1)}_{n,k}=g^{(k-[n+1])}_{n+1,k}-\beta_{n+1}g^{(k-[n+2])}_{n+2,k}$
$=\prod_{s=n+1}^{k-1}\beta_{s}-\beta_{n+1}\prod_{s=n+2}^{k-1}\beta_{s}=0.$ (29)
In its turn, by setting l=nk2l=n-k-2 in Eq. (28) and using the result in Eq. (III.2) for $n+1$ and for $n+2$ we get:
$g^{(k-n-2)}_{n,k}=g^{(k-[n+1]-1)}_{n+1,k}-\beta_{n+1}g^{(k-[n+2]-1)}_{n+2,k}=0.$ (30)
In this way, by decreasing $l$ one step at a time and repeatedly applying Eq. (28), one obtains on the right-hand side only the zero values derived in the preceding steps. This shows that $g^{(l)}_{n,k}=0$ for all l<knl<k-n. Combining this result with the conditions from Eq. (24) we have:
$g^{(l)}_{n,k}=0,\quad l<l_{0}(k,n)\equiv\mathrm{max}(k-n,0)\cup n<0.$ (31)
Now, let us solve the two-dimensional recursion in Eq. (25) under the boundary conditions in Eqs. (26)-(27). To this goal we iterate Eq. (25) with respect to the index “nn” in the decreasing direction, using Eq. (25) itself on the first term on its right-hand side:
$g^{(l)}_{n,k}$ $=$ $g^{(l)}_{n-1,k}+\beta_{n}g^{(l-1)}_{n+1,k}$
$=$ $g^{(l)}_{n-2,k}+\beta_{n-1}g^{(l-1)}_{n,k}+\beta_{n}g^{(l-1)}_{n+1,k}$
$=$ $g^{(l)}_{n-3,k}+\beta_{n-2}g^{(l-1)}_{n-1,k}+\beta_{n-1}g^{(l-1)}_{n,k}+\beta_{n}g^{(l-1)}_{n+1,k}$
$=$ $\ldots.$
Continuing this procedure, the first term eventually reaches the index at which it vanishes by Eq. (31), namely $g^{(l)}_{k-l-1,k}=0$. Hence the iteration yields a one-dimensional recursion in the “ll”-index:
$g^{(l)}_{n,k}=\sum_{s=k-l}^{n}\beta_{s}g^{(l-1)}_{s+1,k}$. (32)
Iteration of the one-dimensional recursion in Eq. (32) produces the nested-sum representation
$g^{(l)}_{n,k}=\sum_{s_{1}=k-l}^{n}\beta_{s_{1}}\sum_{s_{2}=k-l+1}^{s_{1}+1}\beta_{s_{2}}\ldots\sum_{s_{l}=k-1}^{s_{l-1}+1}\beta_{s_{l}},$ (33)
where we have taken into account that $g^{(0)}_{s_{l}+1,k}=1$ by Eq. (26). The expression in Eq. (33) reduces to the previously obtained formula for $k=0$ My1 , in which case the lower limits may be taken as $s_{j}\geq 0$, since $\beta_{p}=0$ for $p<0$.
III.3 The $g$-factors as powers of a Hessenberg matrix
For the special initial state $|\Psi_{0}\rangle$ ($k=0$) the nested sum of Eq. (33) was rewritten in Ref. My1 using some auxiliary Hessenberg matrix. The same approach generalizes to an arbitrary basis state $|\Psi_{k}\rangle$ ($k\geq 0$). Introduce the orthonormal basis of column-vectors
$|e_{0}\rangle\equiv\left(\begin{matrix}1\\ 0\\ \vdots\\ 0\end{matrix}\right),\ldots,|e_{N}\rangle\equiv\left(\begin{matrix}0\\ \vdots\\ 0\\ 1\end{matrix}\right)$ (34)
in an auxiliary linear space of dimension $N+1$, and denote the corresponding row vectors in the dual (conjugate) space by $\langle e_{0}|,\ldots,\langle e_{N}|$. Let matrix $\mathbf{B}$ be as follows
$\mathbf{B}=\sum_{n=0}^{N}|e_{n}\rangle\sum_{s=0}^{n}\beta_{s}\langle e_{s+1}|.$ (35)
The matrix $\mathbf{B}$ of rank $N$ belongs to the class of lower Hessenberg matrices (only the main diagonal and the first superdiagonal may be nonzero). Powers of such matrices are easy to compute numerically Hessen . The $l$th power of $\mathbf{B}$ admits the representation
$\mathbf{B}^{l}=\sum_{n=0}^{N}|e_{n}\rangle\sum_{s_{1}=0}^{n}\beta_{s_{1}}\sum_{s_{2}=0}^{s_{1}+1}\beta_{s_{2}}\ldots\sum_{s_{l}=0}^{s_{l-1}+1}\beta_{s_{l}}\langle e_{s_{l}+1}|.$ (36)
To recover the nested sums of Eq. (33) one evaluates the matrix element
$g^{(l)}_{n,k}=\langle e_{n}|\mathbf{B}^{l}\sum_{m=k}^{N}|e_{m}\rangle.$ (37)
Indeed, Eq. (36) implies $s_{l}\geq k-1$, thus $s_{l-1}\geq s_{l}-1=k-2$ from the upper limit of the innermost nested sum. By utilizing repeatedly the upper limits in Eq. (36), $s_{j-1}\geq s_{j}-1$, we obtain $s_{j}\geq k-l+j-1$, i.e., the lower limits in Eq. (33).
Moreover, introducing the lower triangular matrix $\mathbf{T}$:
$\mathbf{T}=\sum_{n=0}^{N}\sum_{p=0}^{N-n}|e_{n+p}\rangle\langle e_{n}|$ (38)
(the elements in the lower tringular part are all equal to $1$), we get an equivalent expression
$g^{(l)}_{n,k}=\langle e_{n}|\mathbf{B}^{l}\mathbf{T}|e_{k}\rangle.$ (39)
III.4 The state evolution: main results
The above analysis can be summarized as follows.
Theorem 1
The state $|A_{m,k}\rangle\equiv(\hat{A}+\hat{A}^{\dagger})^{m}(\hat{A}^{\dagger})^{k}|\Psi_{0}\rangle$ is given by
$|A_{m,k}\rangle=\sum_{l=0}^{L_{m,k}}g^{(l)}_{k+m-2l,k}(\hat{A}^{\dagger})^{k+m-2l}|\Psi_{0}\rangle,$ (40)
where $L_{m,k}=\mathrm{min}\left(m,\left[\frac{k+m}{2}\right]\right)$ (with $\ldots$] denoting the integer part) and the scalars $g^{(l)}_{n,k}$ are given by Eq. (33) or, alternatively, by Eqs. (37)-(39).
Corollary 1
The time evolution of the rescaled basis state $(-i\hat{A}^{\dagger})^{k}|\Psi_{0}\rangle$ is as follows:
$e^{-i\tau(\hat{A}+\hat{A}^{\dagger})}(-i\hat{A}^{\dagger})^{k}|\Psi_{0}\rangle=\sum_{n=0}^{N}\gamma_{n,k}(\tau)(-i\hat{A}^{\dagger})^{n}|\Psi_{0}\rangle,$
$\gamma_{n,k}(\tau)=\sum_{l=l_{0}(k,n)}^{\infty}\frac{(-1)^{l}\tau^{n-k+2l}}{(n-k+2l)!}g^{(l)}_{n,k}$, (41)
where $l_{0}(k,n)=\mathrm{max}(k-n,0)$. In other words, the quantum amplitude $\gamma_{n,k}(\tau)$ is a power series expansion in $\tau$ with the coefficients $g^{(l)}_{n,k}$ being given by Eq. (33) or, alternatively, by Eqs. (37)-(39).
Proof.– To prove Eq. (1), we substitute the result of Theorem 1 into the power series expansion of the evolution operator, interchange the order of the summation and introduce a new index $n\equiv m+k-2l$:
$e^{-i\tau(\hat{A}+\hat{A}^{\dagger})}(-i\hat{A}^{\dagger})^{k}|\Psi_{0}\rangle=\sum_{m=0}^{\infty}\frac{(-i\tau)^{m}}{m!}(-i)^{k}|A_{m,k}\rangle$
$=\sum_{n=0}^{N}\left(\sum_{l=l_{0}(k,n)}^{\infty}\frac{(-1)^{l}\tau^{n-k+2l}}{(n-k+2l)!}g^{(l)}_{n,k}\right)(-i\hat{A}^{\dagger})^{n}|\Psi_{0}\rangle,$
where we have taken into account that the interval $0\leq l\leq L_{m,k}=L_{n-k+2l,k}$ is equivalent to two conditions: $2l\leq n+2l$ and $l\leq n-k+2l$, the first giving $n\geq 0$, while the second $l\geq\mathrm{max}(0,k-n)=l_{0}(k,n)$. The sum inside the parentheses is the expression for $\gamma_{n,k}(\tau)$ given in Eq. (1). Q.E.D.
Some observations are in order. From the nested-sum expression in Eq. (33) one can infer by induction that $s_{j}\leq n+j-1$, thus:
$g^{(l)}_{n,k}\leq\max_{0\leq p\leq l-1}\left(\sum_{s=k-l+p}^{n+p}\beta_{s}\right)^{l}\leq\left(\sum_{s=k-l}^{n+l-1}\beta_{s}\right)^{l}.$ (42)
Since the sum over $\beta_{n}$ is finite, Eq. (42) ensures that the power series in Eq. (1) converges regardless of the values of the $\beta$-factors (the series of the absolute values is bounded by a uniformly convergent series). Hence, $\gamma_{n,k}(\tau)$ in Eq. (1) is a holomorphic function in the complex plane of $\tau$.
One can give an alternative expression for $\gamma_{n,k}$ as follows:
$\gamma_{n,k}(\tau)$ $=$ $\sum_{l=0}^{\infty}\frac{(-1)^{l+l_{0}(k,n)}\tau^{|n-k|+2l}}{(|n-k|+2l)!}\langle e_{n}|\mathbf{B}^{l+l_{0}(k,n)}\mathbf{T}|e_{k}\rangle$ (43)
$=$ $\sum_{l=0}^{\infty}\frac{(-1)^{l}\tau^{n-k+2l}}{(n-k+2l)!}\langle e_{n}|\mathbf{B}^{l}\mathbf{T}|e_{k}\rangle,$
where in the last sum the actual summation starts with $l\geq l_{0}(k,n)$, since $g^{(l)}_{n,k}$ are all zero for $l<l_{0}(k,n)$ (the matrix average is zero, as can be easily verified). If we introduce a series of matrix functions of $\tau$:
$\mathbf{\Gamma}_{m}\equiv\sum_{l=0}^{\infty}\frac{(-1)^{l}\tau^{m+2l}}{(m+2l)!}\mathbf{B}^{l}\mathbf{T},$ (44)
then (with the above observation on the lower limit of the summation)
$\gamma_{n,k}(\tau)=\langle e_{n}|\mathbf{\Gamma}_{n-k}(\tau)|e_{k}\rangle=\frac{\mathrm{d}^{k}}{\mathrm{d}\tau^{k}}\langle e_{n}|\mathbf{\Gamma}_{n}(\tau)|e_{k}\rangle.$ (45)
The results of Theorem 1 and Corollary 1 reduce to the previously derived results My1 for $k=0$. For instance, it was verified in Ref. My1 by direct substitution and using the recursion Eq. (25) that for $k=0$ Eq. (1) solves the evolution equation. For $k\geq 0$ the verification by direct substitution in Ref. My1 also works since the recursion is the same. The initial condition is guaranteed by the observation on the actual summation below Eq. (43).
From Eqs. (4) and (1)-(43) one immediately obtains the following.
Corollary 2
The average of the evolution operator with the Hamiltonian of Eqs. (2)-(3) on two arbitrary basis states from Eq. (4) reads:
$\langle\Psi_{n}|e^{-i\tau(\hat{A}+\hat{A}^{\dagger})}|\Psi_{k}\rangle=\left[\frac{\prod_{p=0}^{n-1}\beta_{p}}{\prod_{q=0}^{k-1}\beta_{q}}\right]^{\frac{1}{2}}$
$\times\sum_{l=l_{0}(k,n)}^{\infty}\frac{(-i\tau)^{n-k+2l}}{(n-k+2l)!}\langle e_{n}|\mathbf{B}^{l}\mathbf{T}|e_{k}\rangle,$ (46)
where the orthonormal vectors $|e_{m}\rangle$, $m=0,\ldots,N$ span an auxiliary vector space (isomorphic to the considered subspace $\mathcal{H}_{N}$ of the Hilbert space), the lower-Hessenberg matrix $\mathbf{B}$ is defined in Eq. (35) and the lower-triangular matrix $\mathbf{T}$ in Eq. (38). Here $l_{0}(k,n)=\mathrm{max}(0,k-n)$, i.e., for $n<k$ the nonzero terms have $l\geq k-n$. Alternatively, by using Eq. (43), Eq. (2) can be rewritten as
$\langle\Psi_{n}|e^{-i\tau(\hat{A}+\hat{A}^{\dagger})}|\Psi_{k}\rangle=(-i)^{n-k}\left[\frac{\prod_{p=0}^{n-1}\beta_{p}}{\prod_{q=0}^{k-1}\beta_{q}}\right]^{\frac{1}{2}}\gamma_{n,k}(\tau)$
with $\gamma_{n,k}(\tau)$ given by Eq. (1).
IV The energy spectrum
One approach to determining the energy spectrum of a solvable model is to use the algebraic or group-theoretic methods GroupMeth ; LieAlg3Boson ; GroupMeth2 ; LieAlgGen1 ; LieAlgGen2 . Here an alternative approach is developed. The characteristic polynomial for the energy eigenvalues is derived from a recursion, where the coefficients are expressed as nested sums. Alternatively, the characteristic polynomial can be represented by a continued fraction. The amplitudes of the eigenstates are given in several equivalent forms: as continued fractions (or, equivalently, as a product of the Möbius group matrices in two-dimensional projective space), as nested sums, and as principal minors of a Jacobi matrix.
Consider the eigenvalue problem for the Hamiltonian of Eq. (2) in an invariant subspace $\mathcal{H}_{N}$. Expanding eigenstates in the basis of Eq. (4),
$|\Psi(\tau)\rangle=e^{-i\lambda\tau}\sum_{n=0}^{N}\psi_{n}|\Psi_{n}\rangle,$ (48)
we obtain the stationary Schrödinger equation in the following form
$\lambda\psi_{n}=\sqrt{\beta_{n-1}}\psi_{n-1}+\sqrt{\beta_{n}}\psi_{n+1},$ (49)
where $\psi_{-1}=\psi_{N+1}=0$ (recall that $\beta_{-1}=\beta_{N}=0$). Observe also that, for an eigenstate, $\psi_{0}\neq 0$ and $\psi_{N}\neq 0$. Eq. (49) remains invariant under the transformation: $\lambda\to-\lambda$ and $\psi_{n}\to(-1)^{n}\psi_{n}$. Hence, the nonzero eigenvalues come in pairs $\pm\lambda$.
IV.1 The spectrum via continuous fractions
Consider the rescaled ratio of two nearest quantum amplitudes
$R_{n}\equiv\frac{\sqrt{\beta_{n}}\psi_{n}}{\psi_{n+1}},\quad n=0,\ldots,N-1,$ (50)
or its inverse. The boundary conditions on $\psi_{n}$ require that $R_{-1}=0$ and $R_{N}=\infty$. Eq. (49) tells us that the ratio $R_{n}$ satisfies a simple recursion, which can be cast in two equivalent forms:
$R_{n}=\frac{\beta_{n}}{\lambda-R_{n-1}},\quad\frac{\beta_{n}}{R_{n}}=\frac{\beta_{n}}{\lambda-\frac{\beta_{n+1}}{R_{n+1}}}.$ (51)
The boundary conditions are
$\frac{\beta_{0}}{R_{0}}=R_{N-1}=\lambda.$ (52)
One of the relations in Eq. (52) gives $R_{n}$ at either $n=0$ or n=N1n=N-1, while the other – the characteristic equation for the energy eigenvalues $\lambda$. The existence of two equivalent forms for the characteristic equation follows from the finite dimension of the invariant subspace $\mathcal{H}_{N}$ and the inversion symmetry of the row and column indices in the matrix formulation of the eigenvalue problem in $\mathcal{H}_{N}$, i.e., $(\beta_{0},\beta_{1},\ldots,\beta_{N-1})\rightarrow(\beta_{N-1},\beta_{N-2},\ldots,\beta_{0})$ (the characteristic equation is the determinant of a Jacobi matrix, see Eqs. (70)-(71) below).
The solution to Eqs. (51)-(52) can be conveniently cast in the form of two continued fractions:
$R_{n}=\cfrac{\beta_{n}}{\lambda-\cfrac{\beta_{n-1}}{\lambda-\cfrac{\beta_{n-2}}{\cfrac{\ddots}{\lambda-\cfrac{\beta_{0}}{\lambda}}}}},\quad\frac{\beta_{n}}{R_{n}}=\cfrac{\beta_{n}}{\lambda-\cfrac{\beta_{n+1}}{\lambda-\cfrac{\beta_{n+2}}{\cfrac{\ddots}{\lambda-\cfrac{\beta_{N-1}}{\lambda}}}}},\qquad$ (53)
where one of the two boundary conditions in Eq. (52) specifies the lowest denominator in the respective continued fraction and the other – the characteristic equation for $\lambda$.
One can use the Möbius group to represent a continued fraction BookCF . The denominators in the continued fractions of Eq. (53) result from the application of the following Möbius transformation:
$z\to\mathbf{M}_{s}(z)\equiv\frac{-\beta_{s}}{\lambda+z}.$ (54)
Therefore, the continued-fraction representations in Eq. (53) become:
$-R_{n}$ $=$ $\mathbf{M}_{n}\left(\mathbf{M}_{n-1}\left(\ldots\mathbf{M}_{0}(0)\right)\ldots\right),$ (55)
$-\frac{\beta_{n}}{R_{n}}$ $=$ $\mathbf{M}_{n}\left(\mathbf{M}_{n+1}\left(\ldots\mathbf{M}_{N-1}(0)\right)\ldots\right).$ (56)
Introduce the following matrix
$\mathbf{M}_{s}\equiv\left(\begin{matrix}0&-\beta_{s}\\ 1&\lambda\end{matrix}\right).$ (57)
There is a simple relation between the product of matrices $\mathbf{M}_{s}$ and the nested Möbius action of Eqs. (54), (55)-(56). We have
$\left(\begin{matrix}p_{s}\\ q_{s}\end{matrix}\right)\equiv\mathbf{M}_{s}\left(\begin{matrix}cz\\ c\end{matrix}\right),\quad\mathbf{M}_{s}(z)=\frac{p_{s}}{q_{s}},\quad\forall c\neq 0.$ (58)
Hence, instead of the series of nested Möbius transformations, one may alternatively compute the product of the respective Möbius matrices Eq. (57) in Eqs. (55)-(56) and use Eq. (58). For instance,
$R_{n}=-\frac{P_{n}}{Q_{n}},\quad\left(\begin{matrix}P_{n}\\ Q_{n}\end{matrix}\right)\equiv\mathbf{M}_{n}\mathbf{M}_{n-1}\ldots\mathbf{M}_{0}\left(\begin{matrix}0\\ 1\end{matrix}\right).$ (59)
Moreover, the identity
$(1,0)\mathbf{M}_{n}=(0,-\beta_{n})$ (60)
relates the first and the second row in the products of $n$ and n1n-1 Möbius matrices in Eq. (59). Therefore, we can write the matrix product as follows
$\mathbf{M}_{n}\mathbf{M}_{n-1}\ldots\mathbf{M}_{0}=\left(\begin{matrix}-\beta_{n}&0\\ 0&1\end{matrix}\right)\left(\begin{matrix}X_{n}&Y_{n}\\ X_{n+1}&Y_{n+1}\end{matrix}\right),$ (61)
where by setting $n=0$ we get: $X_{0}=0$, $X_{1}=1$ and $Y_{0}=1$, $Y_{1}=\lambda$. Then
$R_{n}=\frac{\beta_{n}Y_{n}}{Y_{n+1}}.$ (62)
Now, in order to find the explicit form of the characteristic polynomial for the energy eigenvalues $\lambda$ in Eq. (52), we will find a recursion for the quantities $Y_{n}$ in Eqs. (61)-(62). Evaluating the product in Eq. (61) by using Eq. (61) for n1n-1 and, alternatively, by multiplying by $\mathbf{M}_{n}$ from the left, we obtain:
$\mathbf{M}_{n}\mathbf{M}_{n-1}\ldots\mathbf{M}_{0}=\left(\begin{matrix}0&-\beta_{n}\\ 1&\lambda\end{matrix}\right)\mathbf{M}_{n-1}\ldots\mathbf{M}_{0}$
$=\left(\begin{matrix}0&-\beta_{n}\\ -\beta_{n-1}&\lambda\end{matrix}\right)\left(\begin{matrix}X_{n-1}&Y_{n-1}\\ X_{n}&Y_{n}\end{matrix}\right)$
$=\left(\begin{matrix}-\beta_{n}&0\\ 0&1\end{matrix}\right)\left(\begin{matrix}X_{n}&Y_{n}\\ X_{n+1}&Y_{n+1}\end{matrix}\right).$
Comparing the matrix entry at position $(2,2)$ on the both sides of the above identity, we get the recursion
$Y_{n+1}=\lambda Y_{n}-\beta_{n-1}Y_{n-1},$ (63)
with the initial values: $Y_{0}=1$ and $Y_{-1}\equiv 0$ ($Y_{-1}$, which appears for $n=0$, is multiplied by $\beta_{-1}=0$). Observe that the characteristic equation $R_{N-1}(\lambda)=\lambda$ Eq. (52) in terms of $Y_{n}$ reads:
$Y_{N+1}(\lambda)\equiv\lambda Y_{N}(\lambda)-\beta_{N-1}Y_{N-1}(\lambda)=0,$ (64)
where we have introduced the quantity $Y_{N+1}$, which satisfies the recursion of Eq. (63), but is absent from the product representation in Eqs. (61)-(62), where we have $n\leq N-1$.
One final observation is in order. From Eqs. (49)-(50) and (62)-(63) one can easily establish the following relation between $\psi_{n}$ and $Y_{n-1}$:
$\psi_{n}\equiv\psi_{0}\left(\prod_{s=0}^{n-1}\beta_{s}\right)^{-\frac{1}{2}}Y_{n}.$ (65)
Indeed, if we use Eq. (65) as an alternative definition of $Y_{n}$, then from Eq. (49) it follows that the quantity $Y_{n}$ satisfies the recursion of Eq. (63) and the initial conditions: $Y_{-1}=0$ and $Y_{0}=1$ (recall that the empty product is equal to $1$).
The characteristic equation for the eigenvalues of Eq. (49) reads: $\psi_{N+1}(\lambda)=0$, where $\psi_{n}(\lambda)$ is the general solution. By Eq. (65) this implies $Y_{N+1}(\lambda)=0$, since the denominator on the right-hand side of Eq. (65) is zero for $n=N+1$ (recall that $\beta_{N}=0$).
IV.2 The characteristic equation for the energy eigenvalues
The recursion in Eq. (63), formally considered for unbounded $n\geq 0$, defines a sequence of polynomials in $\lambda$: $Y_{n}=Y_{n}(\lambda)$. It can be easily solved in a similar was as the recursion in Eq. (28). We have the following result.
Theorem 2
The solution to the recursion in Eq. (63) reads:
$Y_{n}(\lambda)=\sum_{l=0}^{\left[\frac{n}{2}\right]}(-1)^{l}\mathcal{G}^{(l)}_{n-1}\lambda^{n-2l},$ (66)
$\mathcal{G}^{(l)}_{n}\equiv\sum_{s_{1}=0}^{n-1}\beta_{s_{1}}\sum_{s_{2}=s_{1}+2}^{n-1}\beta_{s_{2}}\ldots\sum_{s_{l}=s_{l-1}+2}^{n-1}\beta_{s_{l}},\quad l\geq 1,$
and $\mathcal{G}^{(0)}_{n}\equiv 1$.
Proof.– We will use of the following recursive identity (for $l\geq 1$):
$\mathcal{G}^{(l)}_{n}=\mathcal{G}^{(l)}_{n-1}+\beta_{n-1}\mathcal{G}^{(l-1)}_{n-2}$. (67)
The recursion in Eq. (67) easily follows from another equivalent expression for $\mathcal{G}^{(l)}_{n}$. The index in the $j$th nested sum in Eq. (66) satisfies $2(j-1)\leq s_{j}\leq s_{j+1}-2$, $j=1,\ldots,l-1$. Hence, the nested sums can be arranged in the inverse order, i.e., from $l$ to $1$, as follows
$\mathcal{G}^{(l)}_{n}=\sum_{s_{l}=2(l-1)}^{n-1}\beta_{s_{l}}\sum_{s_{l-1}=2(l-2)}^{s_{l}-2}\beta_{s_{l-1}}\ldots\sum_{s_{1}=0}^{s_{2}-2}\beta_{s_{1}}.$ (68)
Separating the term with $\beta_{n-1}$ in the first sum in Eq. (68),
$\mathcal{G}^{(l)}_{n}=\sum_{s_{l}=2(l-1)}^{n-2}\beta_{s_{l}}\sum_{s_{l-1}=2(l-2)}^{s_{l}-2}\beta_{s_{l-1}}\ldots\sum_{s_{1}=0}^{s_{2}-2}\beta_{s_{1}}$
$+\beta_{n-1}\sum_{s_{l-1}=2(l-2)}^{n-3}\beta_{s_{l-1}}\sum_{s_{l-2}=2(l-3)}^{s_{l-1}-2}\beta_{s_{l-2}}\ldots\sum_{s_{1}=0}^{s_{2}-2}\beta_{s_{1}},$
and replacing the summation index in the last inner sum by $l^{\prime}\equiv l-1$, we arrive at the identity in Eq. (67).
Now let us prove that the polynomial of order $n+1$ on the right hand side of Eq. (66) solves the recursion in Eq. (63). First of all, Eq. (66) reproduces the first two polynomials: $Y_{0}=1$ and $Y_{1}=\lambda$ (for $n=0$ and $n=1$ there is only the term with $l=0$). Now consider the expression for $Y_{n+1}(\lambda)$ provided by Eq. (66). By separating the term with $l=0$ and using the identity (67), we obtain:
$Y_{n+1}(\lambda)=\lambda^{n+1}+\lambda\sum_{l=1}^{\left[\frac{n+1}{2}\right]}(-1)^{l}\mathcal{G}^{(l)}_{n-1}\lambda^{n-2l}$
$+\beta_{n-1}\sum_{l=1}^{\left[\frac{n+1}{2}\right]}(-1)^{l}\mathcal{G}^{(l-1)}_{n-2}\lambda^{n+1-2l}$
$=\lambda\sum_{l=0}^{\left[\frac{n+1}{2}\right]}(-1)^{l}\mathcal{G}^{(l)}_{n-1}\lambda^{(n-1)+1-2l}$
$-\beta_{n-1}\sum_{l^{\prime}=0}^{\left[\frac{n+1}{2}\right]-1}(-1)^{l^{\prime}}\mathcal{G}^{(l^{\prime})}_{n-2}\lambda^{(n-2)+1-2l^{\prime}},$
where we have set $l=l^{\prime}+1$ in the last sum. Let us verify that the upper limits in the above sums conform with Eq. (66), thus giving $Y_{n}(\lambda)$ and $Y_{n-1}(\lambda)$, respectively. In the first sum, by virtue of Eq. (68), the nonzero values of $\mathcal{G}^{(l)}_{n-1}$ have $2(l-1)\leq(n-1)-1$, thus we can replace the upper limit with $l\leq\left[\frac{n}{2}\right]$. By a similar argument, in the second sum we have $2(l-1)\leq(n-2)-1$, i.e., the upper limit can be set to $l\leq\left[\frac{n-1}{2}\right]$. Thus we have arrived at the recursion in Eq. (63). Q.E.D.
In order to obtain the rescaled quantum amplitudes of the energy eigenstates from Eqs. (65)-(66), the parameter $\lambda$ must be set to be one of the energy eigenvalues $\lambda_{j}$, $j=1,\ldots,N+1$, of the stationary Schrödinger equation (49). The characteristic equation is $Y_{N+1}(\lambda)=0$ Eq. (64), with $Y_{N+1}(\lambda)$ given by Eq. (66). We get the following.
Corollary 3
The characteristic polynomial for the energy eigenvalues $\lambda_{j}$, $j=1,\ldots,N+1$ in the invariant subspace $\mathcal{H}_{N}$ reads
$\sum_{l=0}^{\left[\frac{N+1}{2}\right]}(-1)^{l}\mathcal{G}^{(l)}_{N}\lambda^{N+1-2l}=0,$ (69)
where the coefficients $\mathcal{G}^{(l)}_{N}$ are defined by Eqs. (66)-(68).
We have already found that the energy eigenvalues come in pairs $\pm\lambda$, except for the zero eigenvalue. Eq. (69) tells us that the (simple) zero eigenvalue appears in all odd-dimensional invariant subspaces $\mathcal{H}_{N}$, i.e., when $N+1$ is an odd integer.
IV.3 Amplitudes of the eigenstates as Jacobi determinants
There is an alternative expression for the quantum amplitudes $\psi_{n}$ of an eigenstate by the principal minors of the associated Jacobi matrix. Consider the sequence of matrices for $1\leq n\leq N+1$:
$\mathbf{I}_{n}\equiv\sum_{s=0}^{n-1}|e_{s}\rangle\langle e_{s}|,$
$\mathbf{J}_{n}\equiv\sum_{s=0}^{n-2}\sqrt{\beta_{s}}\left(|e_{s+1}\rangle\langle e_{s}|+|e_{s}\rangle\langle e_{s+1}|\right).$
where the vectors $|e_{0}\rangle,\ldots,|e_{N}\rangle$ are defined in Eq. (34). The stationary Schrödinger equation in $\mathcal{H}_{N}$, Eq. (49), is the eigenvalue equation for the Jacobi matrix $\mathbf{J}_{N+1}$, where the energy eigenvalues are the roots of the corresponding characteristic equation, i.e., of the Jacobi determinant
$\mathrm{det}\left(\lambda\mathbf{I}_{N+1}-\mathbf{J}_{N+1}\right)=0.$ (71)
It turns out that the rescaled amplitudes are the principal minors of the above Jacobi determinant:
$Y_{n}=\mathrm{det}\left(\lambda\mathbf{I}_{n}-\mathbf{J}_{n}\right).$ (72)
Indeed, one can verify (by expanding the matrix determinant in Eq. (72) over the last row/column) that the sequence of the determinants in Eq. (72) satisfies the recursion of Eq. (63), where also $Y_{0}=1$ (the null matrix) and $Y_{1}=\lambda$. Therefore, from Eqs. (65) and (72) we obtain:
$\psi_{n}=\psi_{0}\left(\prod_{s=0}^{n-1}\beta_{s}\right)^{-\frac{1}{2}}\mathrm{det}\left(\lambda\mathbf{I}_{n}-\mathbf{J}_{n}\right).$ (73)
IV.4 The stationary state
In the odd-dimensional invariant subspaces $\mathcal{H}_{N}$, i.e, for even $N$, there is the zero eigenvalue $\lambda=0$ of the characteristic polynomial Eq. (69). The corresponding rescaled quantum amplitude $Y_{n}(0)$ can be obtained from Eqs. (66) and (68):
$Y_{2p+1}(0)=0,$ (74)
$Y_{2p}(0)=(-1)^{p}\mathcal{G}^{(p)}_{2p-1}=(-1)^{p}\prod_{s=0}^{p-1}\beta_{2s}.$
The corresponding $\psi_{n}$ of the stationary state in $\mathcal{H}_{N}$ with $N=2M$ can be found from Eq. (65). The (non-normalized) stationary state with $\psi_{0}=1$ reads:
$\psi_{2p+1}=0,\quad 0\leq p\leq M-1,$
$\psi_{2p}=(-1)^{p}\left(\frac{\prod_{s=0}^{p-1}\beta_{2s}}{\prod_{s=0}^{p-1}\beta_{2s+1}}\right)^{\frac{1}{2}},\quad 0\leq p\leq M.$
For example, consider the simplest two-mode bosonic model of Eq. (5) with $m=S=1$ and arbitrary $k$, with the interaction Hamiltonian $\hat{H}=\hat{a}^{\dagger}\hat{b}^{k}+\hat{a}(\hat{b}^{\dagger})^{k}$. In this case from Eq. (10) we obtain
$\beta_{n}=(N-n)\prod_{j=1}^{k}(kn+j).$ (76)
The products in Eq. (IV.4) can be recast in this case as follows:
$\prod_{s=0}^{p-1}\beta_{2s}=2^{p}(M)_{p}\prod_{s=1}^{p}\frac{([2s-1]k)!}{([2s-2]k)!},$
$\prod_{s=0}^{p-1}\beta_{2s+1}=2^{p}\left(M-\frac{1}{2}\right)_{p}\prod_{s=1}^{p}\frac{(2sk)!}{([2s-1]k)!},$ (77)
where we have used the notation $(n)_{p}=n(n-1)\ldots(n-p+1)$ for the the falling factorial. Thus the non-zero amplitudes of the (un-normalized) stationary state read:
$\psi_{2p}=(-1)^{p}\left[\frac{(M)_{p}}{(M-1/2)_{p}}\prod_{s=1}^{p}\frac{\{([2s-1]k)!\}^{2}}{([2s-2]k)!(2sk)!}\right]^{\frac{1}{2}}.$ (78)
Observe that the first factor $\frac{(M)_{p}}{(M-1/2)_{p}}$ in the square root in Eq. (78) grows with $p$, whereas the second factor decreases with $p$. Hence, the absolute values (un-normalized probabilities) $\psi_{2p}^{2}$ have local maxima at $p=0$ (i.e., $n=0$) and $p=M$ ($n=N$), as illustrated in Fig. 1 for $k=2$.
Figure 1: The non-zero squared amplitudes $\psi_{2p}^{2}$ of the stationary state in Eq. (78) vs $p$ for $k=2$ and $N=100$ (the value $\psi_{0}^{2}=1$ was removed for better visibility of the distribution for $p>0$).
Quite similar distribution as in Fig. 1 one obtains for the multi-mode nonlinear boson models. For example, consider the model of Eq. (5) with $m=k_{s}=1$ and $S=2$, i.e., the three-mode bosonic model with the interaction Hamiltonian $\hat{H}=\hat{a}^{\dagger}\hat{b}_{1}\hat{b}_{2}+\hat{a}\hat{b}^{\dagger}_{1}\hat{b}^{\dagger}_{2}$ and
$\beta_{n}=(N-n)(n+1)^{2}.$ (79)
In this case from Eq. (IV.4) one obtains the following (un-normalized) stationary state in the invariant subspace $\mathcal{H}_{2M}$
$\psi_{2p}=(-1)^{p}\left[\frac{(M)_{p}}{(M-1/2)_{p}}\right]^{\frac{1}{2}}\frac{(2p-1)!!}{(2p)!!},$ (80)
where, the first factor (the same as in Eq. (78)) grows with $p$, while the second one decreases with $p$. As the result, we have qualitatively the same behavior of the amplitudes as in Fig. 1.
V Discussion
Exactly solvable models are considered highly useful in applications, provided that the methods leading to the exact solutions are sufficiently simple and the resulting expressions are amenable to further analysis. The usual approach to exactly solvable models, however, relies on rather sophisticated mathematical machinery, such as nonlinear deformations of group algebras GroupMeth ; LieAlg3Boson ; GroupMeth2 ; LieAlgGen1 ; LieAlgGen2 and the Quantum Inverse Scattering Method. QISM .
Exactly solvable boson models are of particular interest not only because of the beautiful mathematical structures they embody but also due to their abundance in quantum physics, e.g., in quantum optics. Perhaps owing to the complicated mathematics involved, the physics literature—quite in parallel with analytical studies in mathematical physics—has primarily relied on numerical simulations and various approximations with poorly defined domains of applicability. Numerical simulations are demanding and resource-intensive, and only quite recently BeyondPA1 ; BeyondPA2 have numerical studies been able to explore the strong-interaction or long-interaction-time regimes now achievable in experiments (e.g., Ref. PumpDep ).
On the other hand, perturbative series expansions involve powers of boson creation and annihilation operators, i.e., operators with unbounded norms, when one adopts the so-called parametric approximation for the strong pump mode. Such perturbative series lack well-defined domains of convergence in the rigorous mathematical sense, which limits their reliability—one is left to verify their validity only by comparison with experimental results. A prominent example of this is the spontaneous down-conversion process in quadratic nonlinear media and its generalizations to higher-order nonlinearities GenSqueez . The standard perturbative approach based on the parametric approximation has been shown to be deficient for higher-order processes, as it leads to a divergence of the system state at a finite interaction time if the pump field is treated not as a separate quantum degree of freedom but merely as a parameter Buzek93 ; Drobny94 . It is, in fact, by chance that no divergence occurs in the quadratic case, where the standard parametric approximation yields the squeezed state of light ReviewSq , which remains regular in norm. Nevertheless, the domain of validity of the standard parametric approximation had long remained undefined. Only very recently has this issue been addressed through direct comparison with a non-perturbative, quasi-exact approach ourPRA . The latter was constructed upon the existence of an exact solution previously derived for the evolution of a special initial state in Ref. My1 .
It was therefore deemed necessary to generalize the exact approach of Ref. My1 to its most general form—the task accomplished in the present work. The previous approach has been extended in this work to the arbitrary state evolution and, additionally, to analysis of the energy spectrum, while relying only on elementary algebra, continued fractions, and the recurrence relations. Remarkably, this framework applies to a broad class of exactly solvable boson models sharing two defining features, as discussed in Section II.
The limitations of the current approach stem from its very generality: since the solution is fully characterized by a single polynomial parameter, $\beta_{n}$ in Eq.(4) of Section II, distinct physical regimes correspond to different choices of this polynomial (see Ref. My1 for details), even though the form of the solution remains the same. Thus, to advance the understanding of specific models, one must go beyond generality and focus on particular choices of $\beta_{n}$ representing distinct physical situations. The significance of the present work lies in providing a rigorous, divergence-free framework for such future analyses of particular models.
VI Conclusion
The analytical solutions to both the state-evolution and energy-spectrum problems have been obtained for a broad class of bosonic models sharing two key features: (i) the interaction Hamiltonian is a sum of two Hermitian-conjugate ladder operators, and (ii) the Hilbert space decomposes into finite-dimensional invariant subspaces under this interaction. Consequently, a single polynomial function of the discrete index labeling the basis states within each invariant subspace fully characterizes the model, governing both its dynamical evolution and spectral properties.
The explicit solution to the arbitrary state evolution problem has been found in the form of an infinite time-series expansion whose coefficients are nested sums involving this polynomial. The characteristic polynomial for the energy eigenvalues has coefficients of a similar structure. The eigenstate amplitudes have been represented equivalently as continued fractions, as products of Möbius-group matrices in the projective space, and as principal minors of the associated Jacobi matrix. The results have important area of applications to the nonlinear quantum-optical models, including $k$-photon down-conversion, which play a central role in modern quantum technologies.
An important open problem is the asymptotic limit as the dimension of the invariant subspace (for instance, the mean photon number in the pump mode) tends to infinity. This problem has important immediate application to the spontaneous down-conversion process, the workhorse of modern quantum optics. In the asymptotic limit one can expect to obtain a simple analytical approximation which would reduce to the common parametric approximation under the appropriate conditions ourPRA . The present work provides a solid foundation for addressing this question in the future.
VII acknowledgements
The author acknowledges financial support, grant number 307507/2023-8, from the National Council for Scientific and Technological Development (Conselho Nacional de Desenvolvimento Científico e Tecnológico) of Brazil.
Appendix A Proof of Eq. (17) of section II
We will use the following expansion (a generalization of the binomial theorem)
$(\hat{A}+\hat{A}^{\dagger})^{m}=\sum_{l=0}^{m}\sum_{P_{\mu}}P_{\mu}\{\hat{A}^{l}(\hat{A}^{\dagger})^{m-l}\},$ (81)
where $P_{\mu}$ is the permutation operator acting on the product of $m$ factors, where $\mu$ runs over different orderings of $m$ factors, with $l$ factors of one type and mlm-l of another type (there are $\binom{m}{l}$ of different orderings). Using Eq. (13) and (14) we get from Eq. (81):
[…]
where the scalar functions $a^{(l)}_{m,k}(n)$ are combinations of $\beta_{n}$ of Eq. (4). Using that $\hat{A}^{p}|\Psi_{0}\rangle=0$ for $p>0$ and introducing the scalars $a^{(l)}_{m,k}\equiv a^{(l)}_{m,k}(0)$, we obtain:
$(\hat{A}+\hat{A}^{\dagger})^{m}(\hat{A}^{\dagger})^{k}|\Psi_{0}\rangle=\sum_{l=0}^{\left[\frac{m}{2}\right]}(\hat{A}^{\dagger})^{k+m-2l}a^{(l)}_{m,k}|\Psi_{0}\rangle$
$+\sum_{l=\left[\frac{m}{2}\right]+1}^{\left[\frac{k+m}{2}\right]}(\hat{A}^{\dagger})^{k+m-2l}a^{(l)}_{m,k}|\Psi_{0}\rangle,$ (83)
where the second sum is non-zero for $2l\leq m+k$. Combining the two sums in Eq. (A) and using that $a^{(l)}_{m,k}(n)$ is nonzero only for $l\leq m$, one arrives arrive at Eq. (17) of section II.

### v2

@@
 Exact analytic solution to the state-evolution problem is presented, applicable to a broad class of solvable bosonic models and arbitrary initial states.
-Moreover, the characteristic equation governing the energy spectrum is derived and the eigenstates are found in the form of continued fractions and as principal minors of the associated Jacobi matrix.
+Moreover, the characteristic equation governing the energy spectrum is derived and the eigenstates are found in the form of continued fractions and as the principal minors of the associated Jacobi matrix.
 The results provide a solid analytical framework for discussion of exactly solvable bosonic models.
@@
 In Section II, the class of bosonic models to which the present approach applies is outlined.
-Section III provides the solution of the state-evolution problem for an arbitrary initial state by evaluating the average of the unitary evolution operator between two arbitrary basis states within each invariant subspace of the Hilbert space.
+Section III provides the solution of the state-evolution problem for an arbitrary initial state by giving the average of the unitary evolution operator between two arbitrary basis states within each invariant subspace of the Hilbert space.
 Theorem1 and Corollaries1 and 2 summarize the main results of this section.
@@
 $\mathbf{B}=\sum_{n=0}^{N}|e_{n}\rangle\sum_{s=0}^{n}\beta_{s}\langle e_{s+1}|.$ (35)
-The matrix $\mathbf{B}$ of rank $N$ belongs to the class of lower Hessenberg matrices (only the main diagonal and the first superdiagonal may be nonzero).
+The matrix $\mathbf{B}$ of rank $N$ belongs to the class of lower Hessenberg matrices (in contrast to the lower triangular matrix, the first super-diagonal can be nonzero also).
 Powers of such matrices are easy to compute numerically Hessen .
@@
 $\mathbf{T}=\sum_{n=0}^{N}\sum_{p=0}^{N-n}|e_{n+p}\rangle\langle e_{n}|$ (38)
-(the elements in the lower tringular part are all equal to $1$), we get an equivalent expression
+(the elements in the lower triangular part are all equal to $1$), we get an equivalent expression
 $g^{(l)}_{n,k}=\langle e_{n}|\mathbf{B}^{l}\mathbf{T}|e_{k}\rangle.$ (39)
@@
 $\times\sum_{l=l_{0}(k,n)}^{\infty}\frac{(-i\tau)^{n-k+2l}}{(n-k+2l)!}\langle e_{n}|\mathbf{B}^{l}\mathbf{T}|e_{k}\rangle,$ (46)
-where the orthonormal vectors $|e_{m}\rangle$, $m=0,\ldots,N$ span an auxiliary vector space (isomorphic to the considered subspace $\mathcal{H}_{N}$ of the Hilbert space), the lower-Hessenberg matrix $\mathbf{B}$ is defined in Eq.
+where the orthonormal vectors $|e_{m}\rangle$, $m=0,\ldots,N$ span an auxiliary vector space (isomorphic to the considered subspace $\mathcal{H}_{N}$ of the Hilbert space), the lower Hessenberg matrix $\mathbf{B}$ is defined in Eq.
 (35) and the lower-triangular matrix $\mathbf{T}$ in Eq.
@@
 Alternatively, the characteristic polynomial can be represented by a continued fraction.
-The amplitudes of the eigenstates are given in several equivalent forms: as continued fractions (or, equivalently, as a product of the Möbius group matrices in two-dimensional projective space), as nested sums, and as principal minors of a Jacobi matrix.
-Consider the eigenvalue problem for the Hamiltonian of Eq. (2) in an invariant subspace $\mathcal{H}_{N}$. Expanding eigenstates in the basis of Eq. (4),
+The amplitudes of the eigenstates are given in several equivalent forms: as continued fractions (or, equivalently, as a product of the Möbius group matrices in two-dimensional projective space), as nested sums, and as the principal minors of a Jacobi matrix.
+Consider the eigenvalue problem in an invariant subspace $\mathcal{H}_{N}$ for the Hamiltonian of Eq. (2). Expanding eigenstates in the basis of Eq. (4),
 $|\Psi(\tau)\rangle=e^{-i\lambda\tau}\sum_{n=0}^{N}\psi_{n}|\Psi_{n}\rangle,$ (48)
@@
 Hence, the nonzero eigenvalues come in pairs $\pm\lambda$.
-IV.1 The spectrum via continuous fractions
+IV.1 The spectrum via continued fractions
 Consider the rescaled ratio of two nearest quantum amplitudes
@@
 (49) reads: $\psi_{N+1}(\lambda)=0$, where $\psi_{n}(\lambda)$ is the general solution.
-By Eq.
-(65) this implies $Y_{N+1}(\lambda)=0$, since the denominator on the right-hand side of Eq.
-(65) is zero for $n=N+1$ (recall that $\beta_{N}=0$).
+We have shown that it also reads $Y_{N+1}(\lambda)=0$ Eq.
+(64).
+Observe that the relation between the two forms by Eq.
+(65) for $n=N+1$ involves the singular first factor on the right-hand side (recall that $\beta_{N}=0$).
 IV.2 The characteristic equation for the energy eigenvalues
-The recursion in Eq.
-(63), formally considered for unbounded $n\geq 0$, defines a sequence of polynomials in $\lambda$: $Y_{n}=Y_{n}(\lambda)$.
-It can be easily solved in a similar was as the recursion in Eq.
-(28).
-We have the following result.
+The recursion in Eq. (63) defines a sequence of polynomials in $\lambda$: $Y_{n}=Y_{n}(\lambda)$. It can be easily solved in a similar way as the recursion in Eq. (28). We have the following result.
 Theorem 2
@@
 and replacing the summation index in the last inner sum by $l^{\prime}\equiv l-1$, we arrive at the identity in Eq. (67).
-Now let us prove that the polynomial of order $n+1$ on the right hand side of Eq.
+Now let us prove that the polynomial of degree $n+1$ on the right hand side of Eq.
 (66) solves the recursion in Eq.
@@
 The corresponding $\psi_{n}$ of the stationary state in $\mathcal{H}_{N}$ with $N=2M$ can be found from Eq. (65). The (non-normalized) stationary state with $\psi_{0}=1$ reads:
-$\psi_{2p+1}=0,\quad 0\leq p\leq M-1,$
+$\psi_{2p+1}=0,\quad 0\leq p\leq M-1,$ (75)
 $\psi_{2p}=(-1)^{p}\left(\frac{\prod_{s=0}^{p-1}\beta_{2s}}{\prod_{s=0}^{p-1}\beta_{2s+1}}\right)^{\frac{1}{2}},\quad 0\leq p\leq M.$
@@
 $\beta_{n}=(N-n)\prod_{j=1}^{k}(kn+j).$ (76)
-The products in Eq. (IV.4) can be recast in this case as follows:
+The products in Eq. (75) can be recast in this case as follows:
 $\prod_{s=0}^{p-1}\beta_{2s}=2^{p}(M)_{p}\prod_{s=1}^{p}\frac{([2s-1]k)!}{([2s-2]k)!},$
@@
 $\beta_{n}=(N-n)(n+1)^{2}.$ (79)
-In this case from Eq. (IV.4) one obtains the following (un-normalized) stationary state in the invariant subspace $\mathcal{H}_{2M}$
+In this case from Eq. (75) one obtains the following (un-normalized) stationary state in the invariant subspace $\mathcal{H}_{2M}$
 $\psi_{2p}=(-1)^{p}\left[\frac{(M)_{p}}{(M-1/2)_{p}}\right]^{\frac{1}{2}}\frac{(2p-1)!!}{(2p)!!},$ (80)
@@
 Nevertheless, the domain of validity of the standard parametric approximation had long remained undefined.
-Only very recently has this issue been addressed through direct comparison with a non-perturbative, quasi-exact approach ourPRA .
+Only very recently has this issue been addressed through direct comparison with a non-perturbative asymptotic approach ourPRA .
 The latter was constructed upon the existence of an exact solution previously derived for the evolution of a special initial state in Ref.
@@
 My1 for details), even though the form of the solution remains the same.
-Thus, to advance the understanding of specific models, one must go beyond generality and focus on particular choices of $\beta_{n}$ representing distinct physical situations.
+Thus, to advance the understanding of specific models, one must go beyond generality and focus on a particular $\beta_{n}$ representing a distinct physical situation.
+Clearly, this task requires development of additional approaches, supplementing the present general framework.
 The significance of the present work lies in providing a rigorous, divergence-free framework for such future analyses of particular models.
@@
 Consequently, a single polynomial function of the discrete index labeling the basis states within each invariant subspace fully characterizes the model, governing both its dynamical evolution and spectral properties.
-The explicit solution to the arbitrary state evolution problem has been found in the form of an infinite time-series expansion whose coefficients are nested sums involving this polynomial.
+The explicit solution to the arbitrary state evolution problem has been found in the form of an infinite time-series expansion whose coefficients are nested sums involving the characterizing polynomial.
 The characteristic polynomial for the energy eigenvalues has coefficients of a similar structure.
-The eigenstate amplitudes have been represented equivalently as continued fractions, as products of Möbius-group matrices in the projective space, and as principal minors of the associated Jacobi matrix.
-The results have important area of applications to the nonlinear quantum-optical models, including $k$-photon down-conversion, which play a central role in modern quantum technologies.
+The eigenstate amplitudes have been represented equivalently as continued fractions, as products of Möbius-group matrices in the projective space, and as the principal minors of the associated Jacobi matrix.
+The results have important area of applications to the nonlinear quantum-optical models, including the $k$-photon down-conversion, which play a central role in modern quantum technologies.
 An important open problem is the asymptotic limit as the dimension of the invariant subspace (for instance, the mean photon number in the pump mode) tends to infinity.
@@
 $(\hat{A}+\hat{A}^{\dagger})^{m}=\sum_{l=0}^{m}\sum_{P_{\mu}}P_{\mu}\{\hat{A}^{l}(\hat{A}^{\dagger})^{m-l}\},$ (81)
-where $P_{\mu}$ is the permutation operator acting on the product of $m$ factors, where $\mu$ runs over different orderings of $m$ factors, with $l$ factors of one type and mlm-l of another type (there are $\binom{m}{l}$ of different orderings).
+where $P_{\mu}$ is the permutation operator acting on the product of $m$ factors, where $\mu$ runs over different orderings of $m$ factors, with $l$ factors of one type and mlm-l of the other type (there are $\binom{m}{l}$ of different orderings).
 Using Eq.